OpenStudy (anonymous):

Find the limit of the function algebraically. lim x->0 (x^2+2x)/(x^4)

hartnn (hartnn):

factor out x from the numerator, what u get ?

OpenStudy (anonymous):

x(x+2)/x^4 ???

hartnn (hartnn):

yes, anything getting cancelled ? simplify that...

OpenStudy (anonymous):

x+2/x^3

hartnn (hartnn):

correct, since you cancelled the factor, that made the numerator = 0 now if you plug in x=0 , numerator is not longer = 0 so just plug in x=0 , what do u get ?

OpenStudy (anonymous):

Error?

OpenStudy (imstuck):

Error is because there is a 0 in the denominator, and mathematically that is illegal!

hartnn (hartnn):

lol, no error, our calculators cannot calculate that anything positive / 0 = \(+\infty \) anything negative / 0 = \(-\infty \) so our numerator is positive or negative ?

OpenStudy (anonymous):

- infinity

hartnn (hartnn):

numerator is negative ? how ?

OpenStudy (anonymous):

2/0?

hartnn (hartnn):

numerator = x+2 since \(x\to 0\), x is very near to 0 so, x+2 will be positive! makes sense ?

hartnn (hartnn):

\(+2/0 =+\infty \)

hartnn (hartnn):

do you have choices/options for this question ?

OpenStudy (anonymous):

lim x->0 (x^2+2x)/(x^4) = x(x+2)/x^4 = x+2/x^3 = lim x-> 0 - (x+2) = 2 = lim x-> 0 - (x^3) = 0 = 2/0 = + infinity

hartnn (hartnn):

that would be correct

OpenStudy (anonymous):

Thank you very much!

hartnn (hartnn):

though more appropriate step would be lim x-> 0 (x+2)/x^3 = (0+2)/0 = +infinity don't break limit into numerator and denominator

OpenStudy (anonymous):

lim x->0 (x^2+2x)/(x^4) = x(x+2)/x^4 = x+2/x^3 = lim x-> 0 (x+2)/x^3 = (0+2)/0 = +infinity